OpenStudy (meilendurcer):

The tenth, eleventh, and twelfth terms of a sequence are shown in the table below: Term number 10 11 12 Term -34 -38 -42 Which of the following shows the first five terms of the sequence? A. -2, 2, 6, 10, 14 B. 2, -2, -6, -10, -14 C. -2, -6, -10, -14, -18 D. 2, 6, 10, 14, 18

OpenStudy (anonymous):

this just looks to be a sequence where each number is added by 4

OpenStudy (anonymous):

you can use the process of elimination by seeing which sequence follows the same pattern.

OpenStudy (meilendurcer):

I don't inderstand the sequence

OpenStudy (anonymous):

And I'm sorry I should have said the sequence is subtracted by 4

OpenStudy (anonymous):

\[-34-4 = -38, -38-4= -42\]

OpenStudy (anonymous):

so you take a look at the answers you are given and see which series follows the same pattern where the numbers become more negative as 4 is subtacted from the previous number within the series

OpenStudy (anonymous):

there are two you can be sure of that are wrong immediately by recognizing this pattern

OpenStudy (anonymous):

let the first number be a and common difference be d [ i said common difference because there's a common difference between the numbers of sequence ] so the 1st term will become a 2nd term will become a+d = d + a = d + first term 3rd term will become a+2d = a +d +d = d+ 2nd term 4th term will become a+3d = a+ d +2d = d+3rd term get it till this point?

OpenStudy (meilendurcer):

Would the closest be C?

OpenStudy (anonymous):

C is correct yes

OpenStudy (anonymous):

well done

OpenStudy (meilendurcer):

yay!!

OpenStudy (anonymous):

do you understand why or was that a guess?

OpenStudy (anonymous):

if you look when you add -4 to -2 you get -6

OpenStudy (meilendurcer):

no i understand. I was also asking to see if i was right.

OpenStudy (anonymous):

ok, then good job

OpenStudy (meilendurcer):

thanks.

OpenStudy (anonymous):

B seems more correct: common difference = -4 so if the first term was 'a', then the tenth term will be \[a - 4*(10 - 1) = -34\] solve this equation for a to get \[a = 2\] so the sequence should start with 2, which means that it can be either B or D Now B is decreasing by 4 whereas D is increasing by 4. Since the given sequence decreases by 4, hence B is correct

OpenStudy (meilendurcer):

I just looked on the internet. Another person said it's B ok thanks for the help.

OpenStudy (anonymous):

its arithmetic sequence, so to get the nth term you have formula nth term = a +(n-1)*d so here 10 th term he gave it as -34 10th term = a + (10-1)*d = a + 9d = -34 \[a + 9*d = -34 ---( 1)\] then 11th term = a + (11 -1)*d = a + 10d = -38 \[a + 10*d = -38 ---(2)\] we know before hand that d is -4, so in (1) if we substitute d value we get a a = -34 -9*(-4) a = -34 + 36 = 2 so first term is 2 , now just -4 ( or subtract 4) to it to get next item and vice versa